A horizontal wire of length 0.27 m, carrying a current of 1.7 A, is placed in a uniform external magnetic field. When the wire is horizontal, it experiences no magnetic force. When the wire is tilted upward at an angle of 19°, it experiences a magnetic force of 1.9 × 10-3 N. Determine the magnitude of the external magnetic field.


1
Expert's answer
2021-03-21T11:37:46-0400
"F_B=BILsin\theta,""B=\dfrac{F_B}{ILsin\theta},""B=\dfrac{1.9\cdot10^{-3}\ N}{1.7\ A\cdot0.27\ m\cdot sin19^{\circ}}=1.27\cdot10^{-2}\ T."